OpenStudy (anonymous):

Without a calculator, determine the square root of 0.00000625 and explain how you figured it out.

OpenStudy (anonymous):

i guess it would help if you knew what the square root of 625 is 25

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i knew that.

OpenStudy (anonymous):

do you just move the decimal over?

OpenStudy (anonymous):

then you would have \[\frac{325}{100,000,000}\]

OpenStudy (saifoo.khan):

take pairs.

hero (hero):

Find two numbers that multiply to get .00000625

OpenStudy (anonymous):

sorry that is \[\frac{625}{100,000,000}\] and the square root of 625 is 25 , the square root of 100,000,000 is 10,000 (half of 8 zeros is 4 zeros) so you get \[\frac{25}{10,000}\] now turn it back into a decimal

OpenStudy (anonymous):

okaaayyy. how do i turn that into a decimal without a calc.....

hero (hero):

.0025 *.0025 This will yield the same thing. It has eight decimal places in total

hero (hero):

I did that without a calculator

OpenStudy (anonymous):

you turn it into a decimal without a calculator by moving the decimal 4 places to the left in the whole number 25 you get hero's answer \[.0025\]

OpenStudy (anonymous):

kk thnx babe

myininaya (myininaya):

nice commas satellite

OpenStudy (anonymous):

thank you. without them i cannot tell 1000000 from 10000000

myininaya (myininaya):

lol

myininaya (myininaya):

whats wrong with \[10^{-8}=\frac{1}{100000000}\]

myininaya (myininaya):

10^(-8) is short and easy to look at it

hero (hero):

myininaya!